A certain weak base has a Kb of 7.10 x 10-7. What concentration of this base will produce a pH of 10.33? concentration:
Answer #1

Sum of pH and pOH is the 14 at 25° c

weak Base Ko=7.10X107 pH = 10-33 PM.+ pol = 14 .pOH = 14-PH = 14-10-33 = 3-67 8+ + op 0 0 = BOH C 7-0 t-t c-ca. por = -log Lo